OpenStudy (anonymous):

Find the general solution of the differential equation 3xdy/dx =2y and then find the value of C if when x=8, y=2. Finally write the particular solution of the equation for the indicated condition.

OpenStudy (bahrom7893):

\[3x\frac{dy}{dx}=2y\]\[3x*dy=2y*dx\]\[\frac{dy}{2y}=\frac{dx}{3x}\]

OpenStudy (bahrom7893):

\[\int\frac{dy}{2y}=\int\frac{dx}{3x}\]\[\frac{1}{2}ln|y|=\frac{1}{3}ln|x|+C\]

OpenStudy (bahrom7893):

sqrt(y)=A*(cube root of x) \[y=A^2*x^{\frac{2}{3}}\]

OpenStudy (anonymous):